# Aligned Probing: Relating Toxic Behavior and Model Internals

### v1

@@ 1 Introduction
We address this gap by introducing aligned probing (Figure 1), a novel interpretability framework (§ 2) that aligns model behavior with internal representations for toxicity. First, we prompt LMs with inputs and assess the toxicity of their generated outputs. Then, during the forward pass, we extract internal representations at each layer to analyze how models encode toxic language. Specifically, we use linear probing (Tenney et al., 2019a; Belinkov, 2022) to train linear models to use these internals to predict specific properties (like input toxicity). Since probes have limited capacity and are rigorously validated (Hewitt and Liang, 2019; Voita and Titov, 2020), their prediction performance on held-out data estimates information strength. Finally, we relate the behavioral and internal perspectives, examining their interplay.
@@ 2.2 Evaluating Model Internals
To evaluate how models encode information about toxicity, we examine the layer-wise information strength of model internals with respect to toxicity scores. We adopt the probing classifier methodology (Tenney et al., 2019a, b; Belinkov, 2022) and approximate information strength () with the performance of a linear model () that maps internal representations () at each layer to toxicity scores ():
Concretely, we first train222For details on training, see Appendix § A.2. a probe to predict from , where the prediction follows:
We then approximate the encoding strength () as the Pearson correlation between the predicted () and actual () toxicity scores. Since the learning capacity of the probe is limited, a high correlation suggests that substantial information about toxicity is encoded in , while a low correlation indicates weaker encoding. Using this method, we formulate four scenarios (Figure 1) to analyze the encoding of input and output toxicity ( and ) within input and output internals ( and ):
@@ 4.2 Internal Evaluation
Figure 4 illustrates how strongly (lines) and consistently (bands) LMs encode the toxicity of text based on the average and standard deviation across 20 probes covering multiple folds and seeds. Our findings challenge previous research that attributes toxicity encoding to upper LM layers (Lee et al., 2024). Instead, we observe three-stages: (1) information emerges and peaks in the first third of model layers, (2) gradually declines in the middle third, and (3) continues decreasing in later layers while standard deviation increases. Notably, the standard deviation (bands) reveals differences even in layers with similar information strength, such as layer one and layer 23, which exhibit deviations of and , respectively. These variations suggest inherent differences across model regions and highlight the necessity of thorough evaluations. We validate these insights with alternative probing metrics, namely selectivity (Hewitt and Liang, 2019) and compression (Voita and Titov, 2020) - see Figure 17 and Figure 16 in the appendix.
@@ Appendix A Appendix
A.1 Limitations
Classifying Toxicity
Detecting toxicity is a non-trivial task as conceptualizations, datasets, and annotator attitudes can vary widely (Waseem, 2016; Waseem et al., 2017; Sap et al., 2022; Pachinger et al., 2023; Cercas Curry et al., 2024). Moreover, toxicity - as with most linguistic properties - is highly contextual and can be implicit, making it difficult to detect (Wen et al., 2023). Even though we consider fine-grained toxicity attributes, our use of PERSPECTIVE API444An industry standard API providing high performance in toxicity detection, see results online. and probing classifiers may miss forms of toxicity not represented in upstream datasets, or exhibit biases (Nogara et al., 2023; Pozzobon et al., 2023).
Beyond English Toxicity
Due to space constraints, we only demonstrate aligned probing with English toxicity in this paper, but the framework is broad; it can be applied to any language model and any textual property. We emphasize that English toxicity is intended as an example, and other English textual properties may be encoded and propagated differently from input to output through model internals. Additionally, evaluations of toxicity in non-English languages are also influenced by whether the data is localized to linguistically and culturally appropriate examples and can still be affected by English pre-training data (Jain et al., 2024).
Probing Classifiers
One of the fundamental problems with using probing classifiers is their limited utility for model explainability. In other words, just because a model’s representations are predictive of a property does not mean that the model is using it (Ravichander et al., 2021; Elazar et al., 2021; Belinkov, 2022). We address this limitation by correlating probing performance with actual toxic behavior when presenting our results, by running causal analyses in addition to correlative analyses, by using control tasks (Hewitt and Liang, 2019) and evaluating our probing setup from an information theory perspective (Voita and Titov, 2020). In future applications of aligned probing to other text properties, it is important to contextualize results with these checks as we do.
A.2 Experimental Details
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]

### v2

@@ 1 Introduction
We address this gap by introducing aligned probing (Figure 1), a novel interpretability framework (§ 2) that aligns model behavior with internals for toxicity. First, we prompt LMs with inputs and assess the toxicity of their generated outputs. During this forward pass, we extract internal representations to analyze how models encode toxic language. Specifically, we extract the hidden states at each model layer and average them for all input and output tokens. Then, we use linear probing (Tenney et al., 2019a; Belinkov, 2022) to estimate the encoded information about the toxicity of the input or output. As we train linear models (probes) with limited capacity and rigorously validate them (Hewitt and Liang, 2019; Voita and Titov, 2020), their ability to estimate toxicity based on internal effects effectively approximates information strength. Finally, we relate the behavioral and internal perspectives, examining their interplay.
@@ 2.2 Evaluating Model Internals
To evaluate how models encode information about toxicity, we extract the hidden states () averaged across all input () and output tokens () at every model layer . We then adopt the probing classifier methodology (Tenney et al., 2019a, b; Belinkov, 2022) to assess how these these internals () linearly map to the corresponding toxicity scores ():
Concretely, we first train222For details on training, see Appendix § A.1. a probe (linear model) to predict from , where the prediction follows:
We then approximate the encoding strength () as the Pearson correlation between the predicted () and actual () toxicity scores. Since the learning capacity of the probe is limited, a high correlation suggests that substantial information about toxicity is encoded in , while a low correlation indicates weaker encoding.
Using this method, we formulate four scenarios (Figure 1) to analyze the encoding of input and output toxicity ( and ) within the averaged input and output internals ( and ):
@@ 4.2 Internal Evaluation
Figure 4 illustrates how strongly (lines) and consistently (bands) LMs encode the toxicity of text based on the average and standard deviation across 20 probes covering multiple folds and seeds. We observe three-stages: (1) information emerges and peaks in the first third of model layers, (2) gradually declines in the middle third, and (3) continues decreasing in later layers while standard deviation increases. Notably, the standard deviation (bands) reveals differences even in layers with similar information strength, such as layer one and layer 23, which exhibit deviations of and , respectively. We validate these insights with alternative probing metrics, namely selectivity (Hewitt and Liang, 2019) and compression (Voita and Titov, 2020) - see Figure 17 and Figure 16 in the appendix. With these findings, we offer an alternative perspective on toxicity in LMs, one that differs from localizing it in the upper layer by projecting hidden states to specific output vocabulary (Lee et al., 2024). Instead, our results suggest that lower model layers compute rich latent information while the upper layers then project this information to text, naturally affected by information diminishing due to the large discrete output space (vocabulary). Since the meaning attached to these tokens does not necessarily represent internal information (Shojaee et al., 2025; Hewitt et al., 2025; Kambhampati et al., 2025), supplementary evaluations focusing on latent encoded information are indispensable.
@@ 9 Discussion and Conclusion
Limitations
Classifying Toxicity
Detecting toxicity is a non-trivial task as conceptualizations, datasets, and annotator attitudes can vary widely (Waseem, 2016; Waseem et al., 2017; Sap et al., 2022; Pachinger et al., 2023; Cercas Curry et al., 2024). Moreover, toxicity - as with most linguistic properties - is highly contextual and can be implicit, making it challenging to detect (Wen et al., 2023). Even though we consider fine-grained toxicity attributes, our use of PERSPECTIVE API444An industry standard API providing high performance in toxicity detection, see results online. and probing classifiers may miss forms of toxicity not represented in upstream datasets, or exhibit biases (Nogara et al., 2023; Pozzobon et al., 2023).
Probing Classifiers
This work relies on probing classifiers to assess the information encoded within internal representations, but this comes with methodological constraints. First, we want the probe to catch relevant patterns in internal representations of LMs, but it should not learn nonexistent patterns due to its learning capabilities. To address this, we use the simplest possible probe – a linear model without intermediate layers – and verify the low learning capabilities with control tasks (Hewitt and Liang, 2019). Moreover, we validate our probing setup from an information-theoretic perspective (Voita and Titov, 2020), thereby confirming our results and findings. Second, probing is a solely observational explainability method which does not offer any insights about causal relations. In other words, just because a model’s representations are predictive of a property does not mean that the model is using it (Ravichander et al., 2021; Elazar et al., 2021; Belinkov, 2022). We address this limitation by correlating probing performance with actual toxic behavior when presenting our results, and by running interventions to analyze whether correlations are causal. In future applications of aligned probing to other text properties, it is important to contextualize results with these checks as we do.
Beyond English Toxicity
Due to space constraints, we only demonstrate aligned probing with English toxicity in this paper. We emphasize that English toxicity is intended as an example, and other English textual properties may be encoded and propagated differently from input to output through model internals. Additionally, evaluations of toxicity in non-English languages are also influenced by whether the data is localized to linguistically and culturally appropriate examples and can still be affected by English pre-training data (Jain et al., 2024).
@@ Appendix A Appendix
A.1 Experimental Details
@@
| OLMo-2 (OLMo et al., 2025) | allenai/OLMo-2-1124-7B | 7 billion | 4.1T tokens |
@@
| Llama-2 (Touvron et al., 2023) | meta-llama/Llama-2-7b-hf | 7 billion | 2T tokens |
@@
| Llama-3 (Grattafiori et al., 2024) | meta-llama/Meta-Llama-3-8B-Instruct | 8 billion | 15T+ tokens |
@@
| Llama-3.1 (Grattafiori et al., 2024) | meta-llama/Llama-3.1-8B | 8 billion | 15T+ tokens |
@@
| Mistral-v0.3 (Jiang et al., 2023) | mistralai/Mistral-7B-v0.3 | 7 billion | unknown |